# Adaptive Transfer Clustering: A Unified Framework

### v2

@@ Acknowledgement
The appendix provides all the proofs of theoretical results in Section 2 and Section 4.
@@ Appendix A Auxiliary lemmas
For any $x\in\mathbb{R}$ and $\Delta>0$,
$\frac{\Delta}{\sqrt{2\pi}}\exp\left(-\frac{\max\left\{x^{2},\left(x+\Delta\right)^{2}\right\}}{2}\right)\leq\left|\Phi\left(x+\Delta\right)-\Phi\left(x\right)\right|\leq\frac{\Delta}{\sqrt{2\pi}}\exp\left(-\frac{\min\left\{x^{2},\left(x+\Delta\right)^{2}\right\}}{2}\right).$
@@ B.1 Proof of Thereom 1
$L(-1,1)\leq\min\{L(1,1),L(1,-1),L(-1,-1)\}\quad\text{or}\quad L(-1,-1)\leq\min\{L(1,1),L(1,-1),L(-1,1)\}$
@@
$\mathbb{P}(\check{Z}^{\lambda}_{0,i}\neq Z^{*}_{0,i})=\mathbb{P}(\check{Z}^{\lambda}_{0,i}\neq Z^{*}_{0,i}|Z^{*}_{0,i}=Z^{*}_{1,i})\mathbb{P}(Z^{*}_{0,i}=Z^{*}_{1,i})+\mathbb{P}(\check{Z}^{\lambda}_{0,i}\neq Z^{*}_{0,i}|Z^{*}_{0,i}\neq Z^{*}_{1,i})\mathbb{P}(Z^{*}_{0,i}\neq Z^{*}_{1,i})$
@@ B.2 Proof of Theorem 2
$\leq\frac{1}{2\sqrt{\pi}\mu}\exp\left(-\mu^{2}\right)+\frac{1}{\sqrt{2\pi}(1+\alpha)\mu}\exp\left(-\frac{\mu^{2}}{2}\left(1+\alpha\right)^{2}\right)+\frac{2}{\sqrt{2\pi}(1-\alpha)\mu}\exp\left(-\frac{\mu^{2}}{2}(1+\alpha)^{2}\right)$
@@
$\leq\frac{1}{2\sqrt{\pi}\mu}\exp\left(-\mu^{2}\right)+\frac{1}{\sqrt{2\pi}(1+\alpha)\mu}\exp\left(-\frac{\mu^{2}}{2}\left(1+\alpha\right)^{2}\right)+\frac{2}{\sqrt{2\pi}(1-\alpha)\mu}\exp\left(-\frac{\mu^{2}}{2}(1+\alpha)^{2}\right)$
@@ B.3 Proof of Theorem 3
$\mathbb{P}_{1,1}\left({\cal E}_{\sf lb}\right)$ $=\mathbb{P}_{1,1}\left(X_{0,i}<-K\right)+\mathbb{P}_{1,1}\left(X_{1,i}<-K\right)+\mathbb{P}_{1,1}\left(X_{0,i}+X_{1,i}<0\right)$
@@
To see it, let ${\cal S}_{0}:=\left\{\left(x,y\right)\in\mathbb{R}^{2}:x<-K,x+y>0\right\}$, ${\cal S}_{1}:=\left\{\left(x,y\right)\in\mathbb{R}^{2}:y>K,x+y<0\right\}$ and $p\left(x\right):=\frac{1}{\sqrt{2\pi}\sigma}\exp\left(-\frac{\left(x-\mu\right)^{2}}{2\sigma^{2}}\right)$. Define $I_{0}:=\int_{{\cal S}_{0}}p(x)p(y)\mathrm{d}x\mathrm{d}y$ and $I_{1}:=\int_{{\cal S}_{1}}p(x)p(y)\mathrm{d}x\mathrm{d}y$. By definition we have $I_{0}=\mathbb{P}_{1,1}\left(\left\{X_{0,i}<-K\right\}\bigcap\left\{X_{0,i}+X_{1,i}>0\right\}\right)$ and $I_{0}+I_{1}=\mathbb{P}_{1,1}\left(X_{0,i}<-K\right)\mathbb{P}_{1,1}\left(X_{1,i}>K\right)$. To show (14), it boils down to show $I_{0}\geq I_{1}$. Notice that
@@ B.4 Proofs of Theorems 4 and 7
(^$\ell\left(\widehat{}\mathbf{Z}_{0}^{\widehat{\lambda}},\bar{}\mathbf{Z}_{0}^{\infty}\right)\leq\widetilde{C}\min_{\lambda\in\Lambda}\left[\Phi\left(-\frac{\sqrt{2}\mu}{\sigma}\right)+\Phi\left(-\frac{\mu}{\sigma}-\frac{\lambda\sigma}{2\mu}\right)\Phi\left(\frac{\mu}{\sigma}-\frac{\lambda\sigma}{2\mu}\right)+\varepsilon\Phi\left(-\frac{\mu}{\sigma}+\frac{\lambda}{2\mu/\sigma}\right)+\frac{\log\left(M/\zeta\right)}{n}\right].$ (18)
@@ B.5 Proof of Theorem 6
We first present the following assumption and result for $d=1$ using the original samples $\left(\mathbf{X}_{0},\mathbf{X}_{1}\right)$ with unknown $\left(\mu,\sigma,\varepsilon\right)$, as the proof for $d\geq 2$ can be reduced to univariate case.
Assumption 4.
for some $\delta_{n}=o(1)$ and $C_{\sf par}\geq 3$.
@@
Now we turn to the proof for $d\geq 2$. Denote $\mu:=\left\|{\bm{\mu}}\right\|$. First note that $\delta_{n}=o\left(\frac{1}{\left(\left\|{\bm{\mu}}\right\|/\sigma\right)^{3}}\wedge\left(\frac{\left\|{\bm{\mu}}\right\|/\sigma}{\log n}\right)^{2}\right)$ implies that $\delta_{n}=o\left(\frac{1}{\left(\widetilde{\mu}/\sigma\right)^{3}}\wedge\left(\frac{\widetilde{\mu}/\sigma}{\log n}\right)^{2}\right)$ by 3, as
@@ C.1 Proof of Corollary 1
$\Phi\left(-\frac{\mu}{\sigma}-\frac{\log\left(\varepsilon^{-1}\left(1-\varepsilon\right)\right)}{2\mu/\sigma}\right)\Phi\left(\frac{\mu}{\sigma}-\frac{\log\left(\varepsilon^{-1}\left(1-\varepsilon\right)\right)}{2\mu/\sigma}\right)\leq\Phi\left(-\frac{2\mu}{\sigma}\left(1+o(1)\right)\right)=o\left(\Phi\left(-\sqrt{2}\mu/{\sigma}\right)\right).$
@@
This concludes the proof.
@@ C.2 Proof of Proposition 1
$\mathbb{P}(\check{Z}^{\textsf{ITL}}_{0,i}\neq Z^{*}_{0,i}\mid Z^{*}_{0,i}=1)=\mathbb{P}\left(X_{0,i}<0\mid Z^{*}_{0,i}=1\right)=\mathbb{P}\bigg{(}\frac{X_{0,i}-\mu}{\sigma}<-\frac{\mu}{\sigma}\bigg{|}Z^{*}_{0,i}=1\bigg{)}=\Phi\left(-\frac{\mu}{\sigma}\right).$
@@
$\mathbb{P}(\check{Z}^{\textsf{DP}}_{0,i}\neq Z^{*}_{0,i})$ $=\mathbb{P}(\check{Z}^{\textsf{DP}}_{0,i}\neq Z^{*}_{0,i}|Z^{*}_{0,i}=Z^{*}_{1,i})\mathbb{P}(Z^{*}_{0,i}=Z^{*}_{1,i})+\mathbb{P}(\check{Z}^{\textsf{DP}}_{0,i}\neq Z^{*}_{0,i}|Z^{*}_{0,i}\neq Z^{*}_{1,i})\mathbb{P}(Z^{*}_{0,i}\neq Z^{*}_{1,i})$
@@ D.1 Proof of Lemma 1
$\mathbb{E}\left(\widehat{\varepsilon}\right)$ $=\mathbb{P}\left(\check{Z}^{\textsf{ITL}}_{0,i}\neq\check{Z}^{\textsf{ITL}}_{1,i}\right)=\varepsilon\cdot\left[\left[1-\Phi\left(-\frac{\mu}{\sigma}\right)\right]^{2}+\left[\Phi\left(-\frac{\mu}{\sigma}\right)\right]^{2}\right]+2\left(1-\varepsilon\right)\Phi\left(-\frac{\mu}{\sigma}\right)\left[1-\Phi\left(-\frac{\mu}{\sigma}\right)\right].$
@@
$\mathbb{E}\left(\frac{\widehat{\varepsilon}}{\varepsilon}-1\right)^{2}$ $\geq\frac{1/\varepsilon}{4n}+\frac{1/\varepsilon^{2}}{4n}\cdot\Phi\left(-\frac{\mu}{\sigma}\right)+\left[\frac{1}{\varepsilon}\cdot\Phi\left(-\frac{\mu}{\sigma}\right)\right]^{2}$
@@ D.2 Proof of Lemma 2
$\leq\left[\widehat{\phi}\left(\widehat{\lambda}\right)+\widehat{\psi}\left(\lambda\right)\right]+\left[\phi\left(\lambda\right)+\widehat{\psi}\left(\lambda\right)+\xi_{\widehat{\psi}}\right]$ (by definition of $(\text{by definition of }\widehat{\phi})$
@@
$\leq\left[\widehat{\phi}\left(\lambda\right)+\widehat{\psi}\left(\lambda\right)\right]+\left[\phi\left(\lambda\right)+2\widehat{\psi}\left(\lambda\right)+\xi_{\widehat{\psi}}\right]$ (by definition of $(\text{by definition of }\widehat{\lambda})$
@@
$\leq\phi\left(\widehat{\lambda}\right)+\widehat{\psi}\left(\widehat{\lambda}\right)+\xi_{\widehat{\psi}}$ (by definition of $(\text{by definition of }\widehat{\phi}\text{~{}and~{}}\xi_{\widehat{\psi}})$
@@
$\leq\phi\left(\lambda\right)+\left[\widehat{\phi}\left(\lambda\right)+\widehat{\psi}\left(\lambda\right)\right]+\xi_{\widehat{\psi}}$ (by definition of $(\text{by definition of }\widehat{\lambda})$

### v3

@@ Acknowledgement
The appendix provides all the proofs of theoretical results.
@@ Appendix A Auxiliary lemmas
For any $x\in\mathbb{R}$ and $\Delta>0$, we have
$\Phi\left(x+\Delta\right)-\Phi\left(x\right)\leq\frac{\Delta}{\sqrt{2\pi}}\exp\left(-\frac{\min\left\{x^{2},\left(x+\Delta\right)^{2}\right\}}{2}\right),$
and
$\Phi\left(x+\Delta\right)-\Phi\left(x\right)\geq\frac{\Delta}{\sqrt{2\pi}}\exp\left(-\frac{\max\left\{x^{2},\left(x+\Delta\right)^{2}\right\}}{2}\right).$
@@
We also present the following assumption and lemma which are needed in our main proofs.
Assumption 4.
for some $\delta_{n}=o(1)$ and $C_{\sf par}\geq 3$.
@@ B.1 Proof of Thereom 1
$L(-1,1)$ $\leq\min\{L(1,1),L(1,-1),L(-1,-1)\}\quad$
$\text{or}\qquad L(-1,-1)$ $\leq\min\{L(1,1),L(1,-1),L(-1,1)\}$
@@
$\mathbb{P}(\check{Z}^{\lambda}_{0,i}\neq Z^{*}_{0,i})$
$=\mathbb{P}(\check{Z}^{\lambda}_{0,i}\neq Z^{*}_{0,i}|Z^{*}_{0,i}=Z^{*}_{1,i})\mathbb{P}(Z^{*}_{0,i}=Z^{*}_{1,i})+\mathbb{P}(\check{Z}^{\lambda}_{0,i}\neq Z^{*}_{0,i}|Z^{*}_{0,i}\neq Z^{*}_{1,i})\mathbb{P}(Z^{*}_{0,i}\neq Z^{*}_{1,i})$
@@ B.2 Proof of Theorem 2
$\leq\frac{1}{2\sqrt{\pi}\mu}\exp\left(-\mu^{2}\right)+\frac{1}{\sqrt{2\pi}(1+\alpha)\mu}\exp\left(-\frac{\mu^{2}}{2}\left(1+\alpha\right)^{2}\right)$
$+\frac{2}{\sqrt{2\pi}(1-\alpha)\mu}\exp\left(-\frac{\mu^{2}}{2}(1+\alpha)^{2}\right)$
@@
$\leq\frac{1}{2\sqrt{\pi}\mu}\exp\left(-\mu^{2}\right)+\frac{1}{\sqrt{2\pi}(1+\alpha)\mu}\exp\left(-\frac{\mu^{2}}{2}\left(1+\alpha\right)^{2}\right)$
$+\frac{2}{\sqrt{2\pi}(1-\alpha)\mu}\exp\left(-\frac{\mu^{2}}{2}(1+\alpha)^{2}\right)$
@@ B.3 Proof of Theorem 3
$\mathbb{P}_{1,1}\left({\cal E}_{\sf lb}\right)$
$=\mathbb{P}_{1,1}\left(X_{0,i}<-K\right)+\mathbb{P}_{1,1}\left(X_{1,i}<-K\right)+\mathbb{P}_{1,1}\left(X_{0,i}+X_{1,i}<0\right)$
@@
To see it, let ${\cal S}_{0}:=\left\{\left(x,y\right)\in\mathbb{R}^{2}:x<-K,x+y>0\right\}$, ${\cal S}_{1}:=\left\{\left(x,y\right)\in\mathbb{R}^{2}:y>K,x+y<0\right\}$ and $p\left(x\right):=\frac{1}{\sqrt{2\pi}\sigma}\exp\left(-\frac{\left(x-\mu\right)^{2}}{2\sigma^{2}}\right)$. Define $I_{0}:=\int_{{\cal S}_{0}}p(x)p(y)\mathrm{d}x\mathrm{d}y$ and $I_{1}:=\int_{{\cal S}_{1}}p(x)p(y)\mathrm{d}x\mathrm{d}y$. By definition we have $I_{0}=\mathbb{P}_{1,1}\left(\left\{X_{0,i}<-K\right\}\bigcap\left\{X_{0,i}+X_{1,i}>0\right\}\right)$ and
$I_{0}+I_{1}=\mathbb{P}_{1,1}\left(X_{0,i}<-K\right)\mathbb{P}_{1,1}\left(X_{1,i}>K\right).$
To show (18), it boils down to show $I_{0}\geq I_{1}$. Notice that
@@ B.4 Proofs of Theorems 4 and 7
$\ell$ (^$\left(\widehat{}\mathbf{Z}_{0}^{\widehat{\lambda}},\bar{}\mathbf{Z}_{0}^{\infty}\right)$ (22)
$\leq\widetilde{C}\min_{\lambda\in\Lambda}\left[\Phi\left(-\frac{\sqrt{2}\mu}{\sigma}\right)+\Phi\left(-\frac{\mu}{\sigma}-\frac{\lambda\sigma}{2\mu}\right)\Phi\left(\frac{\mu}{\sigma}-\frac{\lambda\sigma}{2\mu}\right)+\varepsilon\Phi\left(-\frac{\mu}{\sigma}+\frac{\lambda}{2\mu/\sigma}\right)+\frac{\log\left(M/\zeta\right)}{n}\right].$
@@ B.5 Proof of Theorem 6
We first prove the result for $d=1$ using the original samples $\left(\mathbf{X}_{0},\mathbf{X}_{1}\right)$ with unknown $\left(\mu,\sigma,\varepsilon\right)$ under 4, as the proof for $d\geq 2$ can be reduced to univariate case.
@@
Now we turn to the proof for $d\geq 2$. By definition, for $\forall i\in[n]$ we have
$\left(\widehat{Z}_{0,i}^{\lambda},\widehat{Z}_{1,i}^{\lambda}\right)$ $=\operatorname*{arg\,min}_{u,v\in\{\pm 1\}}\left\{-\left\langle\widehat{\bm{\mu}},u\mathbf{X}_{0,i}+v\mathbf{X}_{1,i}\right\rangle+\lambda\widehat{\sigma}^{2}\mathbb{I}\left(u\neq v\right)\right\}$ (24)
$=\operatorname*{arg\,min}_{u,v\in\{\pm 1\}}\left\{-\left\|\widehat{\bm{\mu}}\right\|\left(\frac{\widehat{\bm{\mu}}^{\top}\mathbf{X}_{0,i}}{\left\|\widehat{\bm{\mu}}\right\|}+\frac{\widehat{\bm{\mu}}^{\top}\mathbf{X}_{1,i}}{\left\|\widehat{\bm{\mu}}\right\|}\right)+\lambda\widehat{\sigma}^{2}\mathbb{I}\left(u\neq v\right)\right\}.$
In view of (24), we can construct projected samples $\big{\{}\widetilde{X}_{0,i},\widetilde{X}_{1,i}\big{\}}_{i=1}^{n}$ using the estimated parameters $\left(\widehat{\bm{\mu}},\widehat{\sigma}\right)$:
$\widetilde{X}_{0,i}:=\frac{\widehat{\bm{\mu}}^{\top}\mathbf{X}_{0,i}}{\left\|\widehat{\bm{\mu}}\right\|},\quad\widetilde{X}_{1,i}:=\frac{\widehat{\bm{\mu}}^{\top}\mathbf{X}_{1,i}}{\left\|\widehat{\bm{\mu}}\right\|},\quad\forall i\in[n].$
We thus have $\textsf{TC}\left(\mathbf{X}_{0},\mathbf{X}_{1};\lambda,{\widehat{\bm{\mu}}},\widehat{\sigma}\right)=\textsf{TC}\big{(}\widetilde{}\mathbf{X}_{0},\widetilde{}\mathbf{X}_{1};\lambda,\left\|\widehat{\bm{\mu}}\right\|,\widehat{\sigma}\big{)}$ for any $\lambda>0$. Note that we can view $\left(\left\|\widehat{\bm{\mu}}\right\|,\widehat{\sigma}\right)$ as fixed parameters, which serves as a good approximation of $\left(\left\|{\bm{\mu}}\right\|,\sigma\right)$ by 3. Denote $\mu:=\left\|{\bm{\mu}}\right\|$ and $\widetilde{\mu}:=\left\|\widehat{\bm{\mu}}\right\|^{-1}{\widehat{\bm{\mu}}^{\top}{\bm{\mu}}}$. Due to independence, we have
$\widetilde{X}_{0,i}|Z_{0,i}^{*}\sim N\big{(}Z_{0,i}^{*}\widetilde{\mu},\sigma^{2}\big{)},\quad\widetilde{X}_{1,i}|Z_{1,i}^{*}\sim N\big{(}Z_{1,i}^{*}\widetilde{\mu},\sigma^{2}\big{)},\quad\forall i\in[n].$
In addition, let
$\widetilde{X}^{b}_{0,i}:=\frac{\widehat{\bm{\mu}}^{\top}\mathbf{X}^{b}_{0,i}}{\left\|\widehat{\bm{\mu}}\right\|},\quad\widetilde{X}_{1,i}:=\frac{\widehat{\bm{\mu}}^{\top}\mathbf{X}^{b}_{1,i}}{\left\|\widehat{\bm{\mu}}\right\|},\quad\forall i\in[n],\quad\forall b\in[B].$
We thus have
$\widetilde{X}_{0,i}^{b}|Z_{0,i}^{*}\sim N\big{(}Z_{0,i}^{*}\left\|\widehat{\bm{\mu}}\right\|,\sigma^{2}\big{)},\quad\widetilde{X}^{b}_{1,i}|Z_{1,i}^{*}\sim N\big{(}Z_{1,i}^{*}\left\|\widehat{\bm{\mu}}\right\|,\sigma^{2}\big{)},\quad\forall i\in[n],\quad\forall b\in[B].$
Similar to (24), we have $\textsf{TC}\left(\mathbf{X}_{0}^{b},\mathbf{X}_{1}^{b};\lambda,{\widehat{\bm{\mu}}},\widehat{\sigma}\right)=\textsf{TC}\big{(}\widetilde{}\mathbf{X}_{0}^{b},\widetilde{}\mathbf{X}_{1}^{b};\lambda,\left\|\widehat{\bm{\mu}}\right\|,\widehat{\sigma}\big{)}$ for any $\lambda>0$. Therefore, ^$\widehat{}\mathbf{Z}_{0}^{\widehat{\lambda}}$ given by ATC using $\big{(}\mathbf{X}_{0},\mathbf{X}_{1}\big{)}$ and $\left(\widehat{\bm{\mu}},\widehat{\sigma}\right)$ is exactly the same as the output of ATC using the projected samples (~$\big{(}\widetilde{}\mathbf{X}_{0},\widetilde{}\mathbf{X}_{1}\big{)}$ and $\left(\left\|\widehat{\bm{\mu}}\right\|,\widehat{\sigma}\right)$, i.e., ^$\widehat{}\mathbf{Z}_{0}^{\widehat{\lambda}}=\textsf{ATC}\left(\mathbf{X}_{0},\mathbf{X}_{1};\Lambda,{\widehat{\bm{\mu}}},\widehat{\sigma}\right)=\textsf{ATC}\big{(}\widetilde{}\mathbf{X}_{0},\widetilde{}\mathbf{X}_{1};\Lambda,\left\|\widehat{\bm{\mu}}\right\|,\widehat{\sigma}\big{)}$. The problem is then reduced to the univariate case with samples $\big{\{}\widetilde{X}_{0,i},\widetilde{X}_{1,i}\big{\}}_{i=1}^{n}$ and unknown $\left(\widetilde{\mu},\sigma,\varepsilon\right)$.
We start by noting that $\delta_{n}=o\left(\frac{1}{\left(\left\|{\bm{\mu}}\right\|/\sigma\right)^{3}}\wedge\left(\frac{\left\|{\bm{\mu}}\right\|/\sigma}{\log n}\right)^{2}\right)$ implies that $\delta_{n}=o\left(\frac{1}{\left(\widetilde{\mu}/\sigma\right)^{3}}\wedge\left(\frac{\widetilde{\mu}/\sigma}{\log n}\right)^{2}\right)$ by 3, as
@@ C.1 Proof of Corollary 1
We first show the following holds:
$\min_{\lambda\in\Lambda}\mathcal{M}\left(\frac{\mu}{\sigma},\varepsilon,\lambda\right)\lesssim\mathcal{M}\left(\frac{\mu}{\sigma},\varepsilon,\log\left(\frac{1-\varepsilon}{\varepsilon}\right)\right).$ (25)
@@
$\Phi\left(-\frac{\mu}{\sigma}-\frac{\log\left(\varepsilon^{-1}\left(1-\varepsilon\right)\right)}{2\mu/\sigma}\right)\Phi\left(\frac{\mu}{\sigma}-\frac{\log\left(\varepsilon^{-1}\left(1-\varepsilon\right)\right)}{2\mu/\sigma}\right)$
$\leq\Phi\left(-\frac{2\mu}{\sigma}\left(1+o(1)\right)\right)=o\left(\Phi\left(-\sqrt{2}\mu/{\sigma}\right)\right).$
@@
This concludes the proof for (25). It suffices to combine the (25), Theorem 4 and Theorem 2 to conclude that the additive term $\log(M\zeta^{-1})/n$ is ignorable provided that $\log\left(\zeta^{-1}\right)=O\left(\log n\right)$ and
$\textsf{SNR}\cdot\min\left\{\left(1+\alpha\right)^{2},2\right\}<\log n.$
@@ C.2 Proof of Proposition 1
$\mathbb{P}(\check{Z}^{\textsf{ITL}}_{0,i}\neq Z^{*}_{0,i}\mid Z^{*}_{0,i}=1)$
$=\mathbb{P}\left(X_{0,i}<0\mid Z^{*}_{0,i}=1\right)=\mathbb{P}\bigg{(}\frac{X_{0,i}-\mu}{\sigma}<-\frac{\mu}{\sigma}\bigg{|}Z^{*}_{0,i}=1\bigg{)}=\Phi\left(-\frac{\mu}{\sigma}\right).$
@@
$\mathbb{P}(\check{Z}^{\textsf{DP}}_{0,i}\neq Z^{*}_{0,i})$
$=\mathbb{P}(\check{Z}^{\textsf{DP}}_{0,i}\neq Z^{*}_{0,i}|Z^{*}_{0,i}=Z^{*}_{1,i})\mathbb{P}(Z^{*}_{0,i}=Z^{*}_{1,i})+\mathbb{P}(\check{Z}^{\textsf{DP}}_{0,i}\neq Z^{*}_{0,i}|Z^{*}_{0,i}\neq Z^{*}_{1,i})\mathbb{P}(Z^{*}_{0,i}\neq Z^{*}_{1,i})$
@@ D.1 Proof of Lemma 1
$\mathbb{E}\left(\widehat{\varepsilon}\right)$ $=\mathbb{P}\left(\check{Z}^{\textsf{ITL}}_{0,i}\neq\check{Z}^{\textsf{ITL}}_{1,i}\right)$
$=\varepsilon\cdot\left[\left[1-\Phi\left(-\frac{\mu}{\sigma}\right)\right]^{2}+\left[\Phi\left(-\frac{\mu}{\sigma}\right)\right]^{2}\right]+2\left(1-\varepsilon\right)\Phi\left(-\frac{\mu}{\sigma}\right)\left[1-\Phi\left(-\frac{\mu}{\sigma}\right)\right].$
@@
$\mathbb{E}\left(\frac{\widehat{\varepsilon}}{\varepsilon}-1\right)^{2}$
$\geq\frac{1/\varepsilon}{4n}+\frac{1/\varepsilon^{2}}{4n}\cdot\Phi\left(-\frac{\mu}{\sigma}\right)+\left[\frac{1}{\varepsilon}\cdot\Phi\left(-\frac{\mu}{\sigma}\right)\right]^{2}$
@@ D.2 Proof of Lemma 2
$\leq\left[\widehat{\phi}\left(\widehat{\lambda}\right)+\widehat{\psi}\left(\lambda\right)\right]+\left[\phi\left(\lambda\right)+\widehat{\psi}\left(\lambda\right)+\xi_{\widehat{\psi}}\right]$ (definition of $(\text{definition of }\widehat{\phi})$
@@
$\leq\left[\widehat{\phi}\left(\lambda\right)+\widehat{\psi}\left(\lambda\right)\right]+\left[\phi\left(\lambda\right)+2\widehat{\psi}\left(\lambda\right)+\xi_{\widehat{\psi}}\right]$ (definition of $(\text{definition of }\widehat{\lambda})$
@@
$\leq\phi\left(\widehat{\lambda}\right)+\widehat{\psi}\left(\widehat{\lambda}\right)+\xi_{\widehat{\psi}}$ (definition of $(\text{definition of }\widehat{\phi}\text{~{}and~{}}\xi_{\widehat{\psi}})$
@@
$\leq\phi\left(\lambda\right)+\left[\widehat{\phi}\left(\lambda\right)+\widehat{\psi}\left(\lambda\right)\right]+\xi_{\widehat{\psi}}$ (definition of $(\text{definition of }\widehat{\lambda})$